# FedEFC: Federated Learning Using Enhanced Forward Correction Against Noisy Labels

### v1

@@ 4.3 Comparison with Baselines
Non-IID variation : To evaluate the impact of non-IID conditions, we fix noise at $(\rho,\zeta)=(0.2,0.8)$ and vary $(\alpha_{\text{dir}},p)$ across three configurations: $(100.0,0.8)$, $(10.0,0.5)$, and $(1.0,0.2)$ (see Table 2). FedEFC achieves the best performance in most configurations. By contrast, FedCorr achieves peak accuracy on MNIST and CIFAR-10 under nearly IID conditions $(100.0,0.8)$ but still lags behind FedEFC and suffers a more significant drop as non-IID characteristics intensify. forward correction performs optimally on MNIST, CIFAR-10, and CIFAR-100 in the most non-IID scenario $(1.0,0.2)$, as it leverages prediction probabilities for all labels, even in their absence from local datasets. However, its accuracy remains closely aligned with FedEFC, differing by at most 0.3% on MNIST and CIFAR-10 and under 1.5% on CIFAR-100. Similarly, confident learning achieves the best accuracy on MNIST under nearly IID $(100.0,0.8)$ and moderately non-IID $(10.0,0.5)$ conditions, yet its performance gap with FedEFC remains within 1%. Despite these cases, FedEFC maintains overall superiority, delivering more stable and resilient performance in non-IID FL environments.
Noise variation : Noise robustness is examined by fixing data allocation and varying noise settings across $(\rho,\zeta)$ pairs: high noise $(0.4,0.8)$, moderate noise $(0.2,0.4)$, and low noise $(0.1,0.0)$, as summarized in Table 3. In most noise conditions, FedEFC demonstrates greater robustness, consistently surpassing the baseline methods, with only minor differences in a few cases. For MNIST under low noise conditions, FedCorr and confident learning marginally outperform FedEFC, with differences of less than 0.3%. Similarly, for MNIST under moderate noise, confident learning attains the highest accuracy, but remains within a 1% margin of FedEFC. Beyond these minor variations, FedEFC consistently yields the best results. Notably, under high noise conditions on CIFAR-10, FedEFC experiences only a 3% performance drop compared to FedAvg in a noise-free setting.
@@ 5 Discussion
Enhanced forward correction : When applying the forward correction in FL, the key difference compared to FedEFC lies in how the noise transition matrix is generated. In FedEFC, the noise transition matrix is derived by counting data that satisfies a model confidence-based threshold, whereas forward correction computes the noise transition matrix directly from the predicted probabilities. In [16], the 97th percentile of the predicted probabilities is used to construct a concrete noise transition matrix. Deriving the noise transition matrix from predicted probabilities does not account for errors in the model predictions, whereas the count-based approach is more resilient to prediction errors, as shown in [15]. In FL with non-IID data, certain labels may be absent from local datasets, leading to reduced accuracy. To mitigate this weakness, class distribution weights are applied to the count matrix, as described in the Implementation Details.
The Nature of the global model in FL : A significant challenge in generating a noise transition matrix is the conventional requirement for a pretrained model. When a model in training is used to construct the noise transition matrix, it tends to overfit to the same dataset used for training the model. For instance, [2] employs an iterative noise cross-validation (INCV) approach that partitions the dataset into separate training and validation set to prevent improper noise transition matrix. In an FL system, the global model obtained prior to local client training is not overfitted to any particular client’s data. As a result, it can effectively generate the noise transition matrix without requiring a pretrained model or data partitioning, unlike conventional methods. FedEFC leverages the inherent properties of FL to overcome this challenge without additional procedures.
@@ 6 Conclusion
We propose FedEFC, a novel algorithm to address the noisy label problem in FL systems with non-IID data distributions. It consists of two phases: (1) finding a prestopping point and (2) applying loss correction. A key advantage of FedEFC is that it effectively mitigates the impact of noisy labels without requiring any inter-client information exchange. In particular, theoretical analysis establishes that, under FedEFC, the FL objective with noisy labels is equivalent to that with clean labels. Extensive experiments across varying noise levels and data heterogeneity demonstrates that FedEFC consistently outperforms conventional FL algorithms. Given that FedEFC operates under realistic conditions with noisy labels and non-IID data, it holds strong potential for adoption in diverse FL scenarios.

### v3

@@ 4.3 Comparison with Baselines
Non-IID variation: To evaluate the impact of non-IID conditions, we fix noise at $(\rho,\zeta)=(0.2,0.8)$ and vary $(\alpha_{\text{dir}},p)$ across three configurations: $(100.0,0.8)$, $(10.0,0.5)$, and $(1.0,0.2)$ (see Table 1). FedEFC consistently achieves the best performance across most settings. On MNIST, RoFL and confident learning achieve slightly higher accuracy under certain non-IID levels, but their gains over FedEFC are within 1%. forward correction performs best under the most severe case $(1.0,0.2)$, yet its accuracy remains comparable to FedEFC, with gaps under 0.3% on MNIST and CIFAR-10, and 1.5% on CIFAR-100. Overall, FedEFC maintains superiority, delivering stable and resilient performance in the non-IID FL condition.
Noise variation: Noise robustness is examined by fixing data allocation and varying noise settings across $(\rho,\zeta)$ pairs: $(0.4,0.8)$, $(0.2,0.4)$, and $(0.1,0.0)$ (see Table 2). In most noise conditions, FedEFC demonstrates greater robustness, steadily surpassing the baseline methods, with only minor differences in a few cases. For the low-noise setting $(0.1,0.0)$ on MNIST, FedCorr and confident learning marginally outperform FedEFC, with differences of less than 0.3%. Under the moderate setting $(0.2,0.4)$ on MNIST, confident learning achieves the highest accuracy, but remains within a 1% margin of FedEFC. In the high-noise setting $(0.4,0.8)$ on CIFAR-10, FedEFC experiences only a 3% performance drop compared to FedAvg in a noise-free setting.
Extreme high noise: To examine robustness under highly adverse conditions, we evaluate CIFAR-10 with extreme noise $(\rho,\zeta)\in\{(0.8,0.0),(0.8,0.4),(0.8,0.8)\}$ (see Table 3). While the standard deviation across trials is higher in this setting, FedEFC remains the most robust method, achieving the highest accuracy in all cases. At $(0.8,0.4)$, the margin over the second-best method FedNoRo is minimal, remaining below $0.5\%$. In contrast, the largest margin occurs at $(0.8,0.8)$, where FedEFC outperforms FedNoRo by more than $2\%$, while maintaining consistent gains across all extreme-noise cases. Detailed configurations and additional results covering all cases of $(\rho,\zeta)$ and $(\alpha_{\text{dir}},p)$ are provided in the supplementary material.
@@ 5 Discussion
Enhanced forward correction: The key distinction between forward correction and FedEFC lies in how the noise transition matrix is constructed. FedEFC derives it by counting samples that satisfy a confidence threshold, whereas forward correction uses predicted probabilities (e.g., 97th percentile in [17]). Unlike probability-based estimation, the count-based approach is more resilient to prediction errors [16]. In FL with non-IID data, where some labels may be absent locally, we further adjust the count matrix using class distribution weights in Sec. 4.2.
Global model in FL: A key challenge in constructing a noise transition matrix is the conventional requirement for a pretrained model, since using a training model often leads to overfitting on the same data. For example, [2] employs iterative noise cross-validation, partitioning the dataset to avoid this issue. In contrast, the global model in FL is not tied to any single client’s data, enabling effective construction of the noise transition matrix without a pretrained model or data partitioning. FedEFC leverages this property to address the challenge without additional procedures.
Clients with heterogeneous noise levels: Another practically relevant scenario arises when different clients are subject to different noise conditions. We evaluated CIFAR-10 under $(\alpha_{\text{dir}},p)=(100.0,0.8)$, where half the clients followed $(\rho,\zeta)=(0.2,0.4)$ and $(\rho,\zeta)=(0.4,0.8)$. In this heterogeneous setup, FedEFC achieved the highest accuracy, demonstrating considerable robustness even when noise patterns varied across clients. Detailed results for this case are provided in the supplementary material.
@@ 6 Conclusion
We propose FedEFC, a novel algorithm to address the noisy label problem in FL systems with non-IID data distributions. It consists of two phases: (1) finding a prestopping point and (2) applying loss correction. A key advantage of FedEFC is that it effectively mitigates the impact of noisy labels without requiring any inter-client information exchange. In particular, theoretical analysis establishes that, under FedEFC, the FL objective with noisy labels is equivalent to that with clean labels. Extensive experiments across varying noise levels and data heterogeneity demonstrates that FedEFC consistently outperforms conventional FL algorithms. Given that FedEFC operates under realistic conditions with noisy labels and non-IID data, it has significant potential for adoption across diverse FL scenarios.
Acknowledgments
This research was partly supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP)-ITRC (Information Technology Research Center) grant funded by the Korea government (MSIT) (IITP-2026-RS-2020-II201787, contribution rate: 50%). Also, this work was supported in part by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00212836, contribution rate: 50%).